OpenStudy (anonymous):

Let f(x)=5x3−60x+5 Input the interval(s) on which f is increasing. (-inf,-2)U(2,inf) nput the interval(s) on which f is decreasing. (-2,2) Find the point(s) at which f achieves a local maximum. Find the point(s) at which f achieves a local minimum. Find the intervals on which f is concave up. Find the intervals on which f is concave down. Find all inflection points.

OpenStudy (anonymous):

I need help with the rest

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

if yall could help solve this really quick, it would mean the world

OpenStudy (triciaal):

did you sketch the function? what is f(0)? did you find the derivative and set = 0 to get the maximum/minimum? find the value of the function at these turning points

OpenStudy (anonymous):

i dont have enough time :(

OpenStudy (anonymous):

i just need the answers

OpenStudy (triciaal):

when x = 2 f(x) = 5(2)^3-60(2) +5 = when x = -2 f(x) = 5(-2)^3-60(-2) + 5= d2y/dx2 = 30x x = 2 +ve so minimum 30(-2) -ve so maximum